### v1

# Comparison of Nested Geometry Treatments within GPU-Based Monte Carlo Neutron Transport Simulations of Fission Reactors

Abstract
Monte Carlo (MC) neutron transport provides detailed estimates of radiological quantities within fission reactors. This method involves tracking individual neutrons through a computational geometry. CPU-based MC codes use multiple polymorphic tracker types with different tracking algorithms to exploit the repeated configurations of reactors, but virtual function calls have high overhead on the GPU. The Shift MC code was modified to support GPU-based tracking with three strategies: (1) dynamic polymorphism (DP) with virtual functions, (2) static polymorphism (SP), and (3) a single tracker (ST) type with tree-based acceleration. Results on the Frontier supercomputer show that the DP, SP, and ST methods achieve 77.8%, 91.2%, and 83.4% of the practical maximum tracking rate in the worst case, indicating that any of these methods can be used without incurring a significant performance penalty. The flexibility of the ST method is highlighted with a hexagonal-grid microreactor problem, performed without hexagonal-grid-specific tracking routines.
1 Introduction
Nuclear reactors account for nearly 20% of electricity production in the United States, with lifecycle greenhouse gas emissions 17–29$\times$ less than coal-fired power plants per unit of energy generated (Center for Sustainable Systems, 2022). These reactors derive energy from induced nuclear fission, a process in which a free neutron is captured by a heavy fuel nucleus, causing it to split into multiple lighter nuclei. Each fission releases heat, as well as additional free neutrons that may induce subsequent fissions, perpetuating the process. The pressurized water reactor (PWR)—the most common reactor design worldwide—consists of a core comprised of assemblies, each of which is comprised of fuel rods, as shown in Figure 1. Water is pumped through the core to extract heat, and electricity is generated with a thermodynamic power cycle via the expansion of steam through a turbine.
Figure 1: Simplified diagram of a generic PWR core shown as a 2D slice. Fuel rods are shown in red and yellow, ostensibly with different fuel compositions.
Significant computational modeling and simulation is required for reactor design, licensing, and operation. Monte Carlo (MC) neutron transport is the preeminent method for obtaining high-fidelity estimates of radiological quantities because of its continuous (i.e., non-discrete) treatment of space, direction, and energy dimensions. This stochastic method involves simulating neutron histories—the circuitous paths that individual neutrons take within a reactor—using a random walk technique. By simulating a large number of histories, accurate statistical estimates of radiological quantities can be deduced. A key radiological quantity is the effective neutron multiplication factor ($k_{\mathrm{eff}}$), defined as the average number of neutrons born from fission that induce subsequent fissions. A $k_{\mathrm{eff}}$ of 1 indicates that the reactor is critical, i.e. operating at steady state, whereas $k_{\mathrm{eff}}<1$ and $k_{\mathrm{eff}}>1$, indicate that the fission rate will decrease or increase over time, respectively. Estimates of $k_{\mathrm{eff}}$ are obtained using MC via a power iteration scheme (Lieberoth, 1968).
Simulating neutron histories requires tracking the positions of neutrons within a computational representation of the reactor, usually a constructive solid geometry (CSG) model. This is accomplished in a fashion similar to the ray tracing techniques used in computer graphics rendering. With CSG, models are constructed from surface primitives (e.g., planes, cylindrical shells, spherical shells) and Boolean logic operations to form cells, i.e., closed regions with uniform material properties, as demonstrated in Figure 2. Tracking operations such as those summarized in Table 1 are implemented by querying the surfaces that comprise each cell.
Figure 2: Demonstration of the CSG construction process, shown as a 2D slice. Plane surfaces (orange), have surface normals (yellow), denoting which side of the plane is positive. The outside of the blue cylindrical surface is considered positive. Cell construction logic is used to form cells 1 and 2. The cell 2 logic can be read as “the intersection of the space on the positive side of surface 1, the negative side of surface 2, the positive side of surface 3, … etc.”
Table 1: High-level geometry tracking operations required by MC transport codes.
Tracking operation Inputs Output
find_cell (1) position (1) cell containing the position
distance_to_surface (1) position (1) distance to the next surface
(2) direction (2) next surface
move_within_cell (1) position (1) new position
(2) direction
(3) distance
cross_surface (1) position (1) new cell after crossing the surface
(2) current cell
(3) current surface
change_direction (1) new direction -
The nested and repeated structures that comprise reactors receive special treatment for performance and user convenience. A CSG universe is a contiguous geometric region comprised of one or more cells that can be embedded within one or more parent cells (West et al., 1979). Array universes are special universe types in which cells comprise a structured mesh. The cells in an array universe must be filled with other universes which may be either CSG universes or other array universes. Examples of CSG universes embedded in rectilinear and hexagonal array universes are shown in Figure 3. Some fission reactors, including PWRs, can be modeled with three levels of universes: an array universe representing the core, with each core array cell filled with an assembly array universe, and each assembly array cell filled with a CSG universe containing fuel rods surrounded by coolant (referred to as a pin within this work).
Figure 3: Demonstration of CSG universes nested within rectilinear and hexagonal arrays universes, shown as a 2D slice.
The use of array universes in concert with CSG universes provides tracking performance benefits. When determining what cell contains a given point, the hierarchical configuration of nested universes can be exploited by first finding which assembly contains the point, then which pin within the assembly, and finally which cell within the pin. This is loosely analogous to standard ray tracing acceleration structures such as bounding volume hierarchy (BVH) (Ericson, 2004) or $k$-d trees (Bentley, 1975). In addition, because CSG and array universes represent geometry differently, the tracking operations summarized in Table 1 are implemented separately for each universe type, allowing each to be optimized. Within this work, objects that carry out these tracking operations are referred to as trackers. Thus, a CSG universe has a corresponding CSG tracker, a rectilinear array universe has a rectilinear array tracker, etc. Array trackers can implement tracking operations significantly more efficiently than CSG trackers by exploiting the regular structure of array universes.
Using multiple tracker types requires tracking functions to be called polymorphically. On the CPU with C++, this is accomplished by implementing tracker functions as virtual functions of a tracker base class. However, using GPU programming models (e.g., HIP, CUDA), it is not clear whether the performance benefits of different tracker types outweigh the overhead associated with virtual function calls (Zhang et al., 2021).
Shift is a general purpose MC radiation transport code developed at Oak Ridge National Laboratory (ORNL) capable of simulating the behavior of neutrons and high-energy photons for fission, fusion, and national security applications (Pandya et al., 2016). Shift, written in C++ with abstracted HIP/CUDA device programming models, supports both CPU and GPU execution (Hamilton and Evans, 2019) and is designed to scale effectively from laptops to leadership-class supercomputers. On the CPU, Shift uses the Oak Ridge Adaptable Nested Geometry Engine (ORANGE) (Johnson et al., 2023) [^1] The ORANGE package is shared between Shift and the Celeritas MC code, used for high-energy physics detector analysis (Johnson et al., 2021)., which includes CSG, rectilinear array, and hexagonal array trackers, leveraged via virtual functions. On the GPU, Shift—as well other MC codes that support GPU execution such as OpenMC (Romano et al., 2015; Tramm et al., 2022), and PRAGMA (Choi et al., 2019)—have avoided GPU polymorphism by supporting only a single universe/tracker type.
Prior to this work, Shift only supported the Reactor Tool Kit (RTK) universe type on the GPU. RTK is a special-purpose universe type that models a full reactor core, and is separate from the ORANGE package. A single RTK universe always consists of (1) a rectilinear array representing the core, (2) rectilinear arrays representing assemblies, and (3) cuboids containing concentric cylinders, representing pins. The RTK tracker is optimized to track neutrons through this specific configuration. As a result, the RTK tracker is expected to provide the best possible performance. However, RTK has limited applicability because not all fission reactors can be modeled with nested rectilinear grids. Numerous reactor designs consist of hexagonal assemblies (Habush and Harris, 1968; Hejzlar et al., 2013; Betzler et al., 2020), and pebble bed reactors consist of irregular configurations of spherical fuel elements (Andreades et al., 2016; Mulder and Boyes, 2020). In addition, MC neutron and photon transport is used for other applications such as nuclear fusion reactors (Juarez et al., 2021; Kos et al., 2023) and accelerator devices (Radel and Van Abel, 2016; Nelson et al., 2022), which involve geometries with complexity far beyond rectilinear grids.
In this work, three methods for GPU-based multi-universe tracking were implemented in ORANGE to assess the trade-offs between single and multiple tracker types. The timing results for Shift $k_{\mathrm{eff}}$ calculations on the Summit (Oak Ridge Leadership Computing Facility, 2023b) and Frontier (Oak Ridge Leadership Computing Facility, 2023a) supercomputers at ORNL were recorded for each of the three methods and compared to RTK timing results. The three methods are outlined as follows. The dynamic polymorphism (DP) method employs multiples tracker types with virtual functions. The static polymorphism (SP) method employs multiple tracker types with switch statements. The single tracker (ST) method avoids polymorphism by using a single tracker type for all universe types. The ST method is accomplished using pseudo-array universes, which are arrays modeled as CSG universes and tracked upon with a CSG tracker. Although this approach forfeits the aforementioned benefits of array trackers, this is counteracted with a bounding interval hierarchy (BIH) acceleration structure (Wächter and Keller, 2006).
The remainder of this work proceeds as follows. Section 2 provides background on how Shift solves for $k_{\mathrm{eff}}$, including the role of tracking operations. Section 3 introduces tracking algorithms for CSG and array universes, multi-universe geometries, and RTK universes. Section 4 describes the implementation of each of the three experimental multi-universe GPU tracking methods. Section 5 describes the computer hardware on Summit and Frontier used for this work. Section 6 provides a performance comparison of the three methods using a model of the NuScale small module reactor (SMR) (NuScale Power, 2023), comprised of rectilinear assemblies. Section 7 demonstrates the flexibility of the ST method on the Empire microreactor benchmark problem (Lee et al., 2020; Matthews et al., 2021) comprised of hexagonal assemblies.
2 Background
This section provides background on how Shift solves for $k_{\mathrm{eff}}$. Section 2.1 describes the neutron transport equation formulated with $k_{\mathrm{eff}}$ as the dominant eigenvalue. Section 2.2 describes the power iteration solution method and the role of tracking operations in the MC random walk process. Section 2.3 describes the Shift implementation of MC power iteration on the GPU.
2.1 Neutron transport equation
Neutrons interact with fuel and non-fuel nuclei via a variety of different mechanisms which are broadly organized into three categories:
  1. 1.
    absorption: a neutron is captured by a nucleus,
  2. 2.
    fission: a neutron is captured by a nucleus, which subsequently breaks apart,
  3. 3.
    scattering: a neutron interacts with a nucleus, effectively changing the neutron’s kinetic energy and direction.
As a neutron travels through a material, the probability that it will undergo one of these interactions is quantified using nuclear cross sections, functions which depend on the energy of the neutron, the types of nuclei within the material, and the density and temperature of the material. Absorption and fission cross sections tend to be negatively correlated with neutron energy. As a result, within a reactor core, neutrons typically scatter $\sim 10^{1}$ times before being terminated via absorption/fission. These effects are quantified in the neutron transport equation:
$(\hat{\mathrm{T}}-\hat{\mathrm{S}})\psi=\frac{1}{k_{\mathrm{eff}}}{\hat{\mathrm{\chi}}}\hat{\mathrm{F}}\psi\>,$ (1)
where $\hat{\mathrm{T}}$ is the transport operator, $\hat{\mathrm{S}}$ is the scattering operator, $\hat{\mathrm{\chi}}$ is energy spectrum of neutrons born from fission, $\hat{\mathrm{F}}$ is the fission operator, and $\psi$ is the neutron flux, which describes the number of neutrons that pass through a 2D area per unit time per unit angle. The $\hat{\mathrm{T}}$, $\hat{\mathrm{S}}$, and $\hat{\mathrm{F}}$ operators depend on nuclear cross sections, but these relationships are omitted for brevity. By collecting the operators in a single term,
$\hat{\mathrm{A}}\equiv\hat{\mathrm{F}}(\hat{\mathrm{T}}-\hat{\mathrm{S}})^{-1}{\hat{\mathrm{\chi}}}\>,$ (2)
Equation 1 can be formulated as a standard eigenvalue problem,
$k_{\mathrm{eff}}\text{f}=\hat{\mathrm{A}}\text{f},$ (3)
where $k_{\mathrm{eff}}$ is the dominant eigenvalue of $\hat{\mathrm{A}}$ and f, the eigenvector, is the fission source given by
$\text{f}=\hat{\mathrm{F}}\psi.$ (4)
Physically, f is a probability density function describing the spatial distribution of fission neutrons. Like $k_{\mathrm{eff}}$, this distribution is not known a priori. Due to the complexity of the geometry and cross sections, Equation 3 must be solved numerically for most practical cases.
2.2 Monte Carlo power iteration
The eigenvalue problem presented in Equation 3 can be solved with MC power iteration. With this method, $k_{\mathrm{eff}}$ and f are solved iteratively by applying the standard power iteration eigenvalue algorithm (von Mises and Pollaczek-Geiringer, 1929) to Equation 3,
$\text{f}^{(n+1)}=\frac{1}{k_{\mathrm{eff}}^{(n)}}\hat{\mathrm{A}}\text{f}^{(n)},$ (5)
$k_{\mathrm{eff}}^{(n+1)}=k_{\mathrm{eff}}^{(n)}\frac{\langle\text{f}^{(n+1)}\rangle}{\langle\text{f}^{(n)}\rangle},$ (6)
[…]
Algorithm 1 MC power iteration algorithm for calculating the converged $k_{\mathrm{eff}}$ using $m_{i}$ inactive cycles, $m_{a}$ active cycles, and $n$ histories per cycle.
1procedure mc_power_iteration($m_{i}$, $m_{a}$, $n$)
2     Set f to initial guess
3     for inactive_cycle $\in[1,m_{i}]$
4         $k_{\mathrm{eff}}$, f $=$ run_histories(f, $n$)
5     for active_cycle $\in[1,m_{a}]$
6         $k_{\mathrm{eff}}$, f $=$ run_histories(f, $n$)
7         accumulate $k_{\mathrm{eff}}$
8     return average of all stored $k_{\mathrm{eff}}$
The run_histories MC transport function used in Algorithm 1 is shown in Algorithm 2. For each of $n$ histories, a neutron birth position is sampled from the supplied guess for the fission source distribution. A random walk technique is then used to move the neutron through a computational representation of the reactor. This algorithm makes use of the geometry tracking operations defined in Table 1. Tracking operations typically account for 20% or less of the total runtime within an MC simulation (Hamilton and Evans, 2019). Cross section calculations make up the plurality of runtime because of the large number of memory fetch operations. By definition, the random walk process imposes divergent neutron paths across histories, resulting in random memory access. As a result, MC simulations tend to be latency bound.
Algorithm 2 MC neutron transport algorithm for simulating $n$ neutron histories born from f, in order to update estimates of $k_{\mathrm{eff}}$ and f. This algorithm uses the geometric tracking operations specified in Table 1.
1procedure run_histories(f, $n$)
2     for history $\in[1,n]$
3         sample position ($r$) from f
4         sample energy ($E$) from fission energy spectrum
5         sample direction ($\Omega$) isotropically
6         cell ($c$) = find_cell ($r$)
7         while true
8              calculate the cross section ($\Sigma$) in $c$
9              distance ($d$), surface ($s$) = distance_to_surface ($r$, $\Omega$)
10              sample ##\## of mean free paths ($\tau$) before event
11              while $d<\tau/\Sigma$
12                  $\tau$ $=$ $\tau$ $\Sigma$$\times$ d
13                  $r$ = move_within_cell ($r$, $\Omega$, $d$)
14                  $c$ = cross_surface ($r$, $c$, $s$)
15                  calculate $\Sigma$ in $c$
16                  $d$, $s$ = distance_to_surface ($r$, $\Omega$)
17              $r$ = move_within_cell ($r$, $\Omega$, $\tau/\Sigma$)
18              sample event type ($\xi$)
19              if $\xi$ == scatter
20                  update $E$ and $\Omega$
21              else if $\xi$ == fission
22                  calculate and store an estimate of $k_{\mathrm{eff}}$
23                  store fission sites
24                  break
25              else if $\xi$ == absorption
26                  break
27     calculate updated $k_{\mathrm{eff}}$ from stored $k_{\mathrm{eff}}$ values
28     calculate updated f from stored fission sites
29     return updated $k_{\mathrm{eff}}$, updated f
2.3 GPU-based implementation in Shift
On the CPU, Shift carries out MC power iteration using the MC neutron transport simulation algorithm shown in Algorithm 2. On the GPU, Shift uses an event-based MC algorithm. As in Algorithm 2, the event-based algorithm involves simulating $n$ histories per cycle; however, operations are reordered to take advantage of single instruction, multiple threads (SIMT) parallelism. All operations—including birth, tracking operations, and collisions—are performed on a vector of histories. This vector is masked in order to only perform operations on applicable histories. Each of the tracking operations in Table 1 is called for a vector of histories via a kernel of the same name. Full details of the event-based transport algorithm are found in Hamilton and Evans (2019). This approach leads to smaller kernel sizes and therefore increased occupancy and higher tracking rates.
3 Tracking algorithms
This section discusses the implementation of the Table 1 tracking operations for different universe types. For brevity, only the two most algorithmically interesting tracking operations are discussed: find_cell and cross_surface. Section 3.1 provides possible implementations of these operations for CSG and rectilinear array universes, where the latter are shown to have significantly lower time complexity. Hexagonal array universe tracking algorithms can be implemented using the same strategy as that used for rectilinear arrays, with the same time complexity. However, hexagonal array indexing is considerably more complicated, so these algorithms are omitted here for simplicity. Section 3.2 provides the implementation of multi-universe tracking on the CPU within ORANGE. Section 3.3 provides the tracking algorithms for the reactor-specific RTK universe type, used by Shift on the GPU.
3.1 Single universe tracking
Possible implementations of find_cell and cross_surface for a CSG tracker are shown in Algorithms 3 and 4. Because CSG universes have no underlying structure, find_cell involves conducting an $O(N)$ search over all of the cells within the universe. The cross_surface implementation relies on a list of neighbor cells connected to each surface, which can be generated as a preprocessing step. This neighbor list can then be searched in $O(N)$ time.
Algorithm 3 CSG tracker version of find_cell.
1procedure find_cell(pos)
2     for cell $\in$ cells
3         if cell contains pos
4              return cell
Algorithm 4 CSG tracker version of cross_surface.
1procedure cross_surface(cell, surf, pos)
2     for new_cell $\in$ neighbors(surf)
3         if new_cell == cell
4              continue
5         if new_cell contains pos
6              return new_cell
Possible implementations of find_cell and cross_surface for a rectilinear array tracker are shown Algorithms 5 and 6. These algorithms have favorable time complexity relative to their CSG counterparts. Assuming the rectilinear grid is $N^{1/3}\times N^{1/3}\times N^{1/3}$, find_cell can be performed in $O(\log(N^{1/3}))$ time by conducting a binary search over the mesh divisions in each dimension. The cross_surface operation can be performed in $O(1)$ time because the surface of each cell is known to have only one neighbor. It requires the current cell, the $i/j/k$ dimension of the surface being crossed, and whether the surface is being crossed in the positive or negative direction.
Algorithm 5 Rectilinear array tracker version of find_cell.
1procedure find_cell(pos)
2     for dim $\in[i,j,k]$
3         ijk[dim] = binary_search(mesh[dim], pos[dim])
4     cell = ijk_to_cell(ijk)
5     return cell
Algorithm 6 Rectilinear array tracker version of cross_surface.
1procedure cross_surface(cell, dim, dir_sign)
2     ijk = cell_to_ijk(cell)
3     if dir_sign is positive
4         ijk[dim] = ijk[dim] + 1
5     else
6         ijk[dim] = ijk[dim] - 1
7     next_cell = ijk_to_cell(ijk)
8     return next_cell
3.2 Multi-universe CPU tracking in ORANGE
The algorithms shown in Section 3.1 show how CSG and array trackers operate within single universes. Here, tracking algorithms for geometries consisting of multiple nested universes are shown, as implemented on the CPU in ORANGE. The ORANGE CPU version of find_cell and cross_surface are shown in Algorithms 7 and 8. The find_cell algorithm recursively finds the cell within daughter universes until reaching the bottom-most (i.e., most embedded) level. In cross_surface, a surf_universe argument is supplied that denotes the top-most (i.e., least embedded) level for which the neutron is on a surface, noting that it may be on a surface on multiple levels. The cell on the other side of the surface is then found by recursing through daughters starting at this level. Both algorithms assume the existence of a polymorphic get_tracker function which returns a tracker object for a given universe, depending on its type.
Algorithm 7 Multi-universe version of find_cell.
1procedure find_cell(pos)
2     tracker $=$ get_tracker(root_universe)
3     cell $=$ tracker.find_cell(pos)
4     while cell.daughter
5         tracker $=$ get_tracker(cell.daughter)
6         cell $=$ tracker.find_cell(pos)
7     return cell
Algorithm 8 Multi-universe version of cross_surface. For simplicity, args can be assumed to be a struct containing the union of the arguments to the standard CSG and array versions of cross_surface.
1procedure cross_surface(surf_universe, args)
2     tracker $=$ get_tracker(surf_universe)
3     next_cell $=$ tracker.cross_surface(args)
4     while next_cell.daughter
5         tracker $=$ get_tracker(daughter)
6         next_cell $=$ tracker.find_cell(args.pos)
7         daughter $=$ cell.daughter
8     return next_cell
3.3 RTK tracking
As mentioned in Section 1, RTK is the reactor-specific universe type currently employed by Shift for GPU execution. RTK is a template universe type explicitly instantiated to contain three nested levels of universes necessary to model reactors such as PWRs with rectilinear configurations. An RTK universe consists of a rectilinear array core universe populated with rectilinear array assembly universes, each populated with pin universes. The pin universe is not a full, general purpose CSG universe, but rather a limited CSG universe consisting of concentric cylinders within a rectangular cuboid cell.
The RTK version of find_cell and cross_surface are shown in Algorithms 9 and 10. These algorithms resemble the standard multi-universe tracking algorithms shown in in Section 3.2, with two key differences. First, because there are exactly three levels of universes, the while loops over daughter universes can be unrolled (in practice, this is achieved through C++ template recursion). Second, the types of universes at each level are known at compile time. Whereas the multi-universe tracking algorithms shown in Section 3.2 rely on a polymorphic get_tracker function, Algorithms 9 and 10 can call non-polymorphic get_rect_tracker and get_csg_tracker functions. As a result of these simplifications, RTK is expected to be the most performant tracker type. However, the clear limitations on geometric complexity imposed by RTK motivate the implementation of a general-purpose GPU tracker type.
Algorithm 9 RTK version of find_cell.
1procedure find_cell(pos)
2     core_tracker $=$ get_rect_tracker(core_universe)
3     core_cell $=$ core_tracker.find_cell(pos)
4     assm_universe = core_cell.daughter
5     assm_tracker $=$ get_rect_tracker(assm_universe)
6     assm_cell $=$ assm_tracker.find_cell(pos)
7     pin_universe = assm_cell.daughter
8     pin_tracker $=$ get_csg_tracker(pin_universe)
9     cell $=$ pin_tracker.find_cell(pos)
10     return cell
Algorithm 10 RTK version of cross_surface.
1procedure cross_surface(surf_universe, args)
2     if surf_universe.level == core
3         core_tracker $=$ get_rect_tracker(surf_universe)
4         core_cell $=$ core_tracker.cross_surface(args)
5         assm_universe $=$ core_cell.daughter
6         assm_tracker
7                $=$ get_rect_tracker(assm_universe)
8         assm_cell $=$ assm_tracker.find_cell(args.pos)
9         pin_universe = assm_cell.daughter
10         pin_tracker $=$ get_csg_tracker(pin_universe)
11         new_cell $=$ pin_tracker.find_cell(args.pos)
12     else if surface_universe.level == assembly
13         assm_tracker $=$ get_arr_tracker(surf_universe)
14         assm_cell $=$ assm_tracker.find_cell(args.pos)
15         pin_universe = assm_cell.daughter
16         pin_tracker $=$ get_csg_tracker(pin_universe)
17         new_cell $=$ pin_tracker.find_cell(args.pos)
18     else if surface_universe.level = pin
19         pin_tracker $=$ get_csg_tracker(surf_universe)
20         new_cell $=$ pin_tracker.find_cell(args.pos)
21     return new_cell
4 Methodology: Multi-universe GPU tracking methods
This section describes the ORANGE implementation of the three experimental GPU-based multi-universe tracking methods explored in this work. The DP and SP methods, which rely on universe-specific tracker types, were implemented only for CSG and rectilinear array universes for the purposes of this work. The ST method, which relies on only a single tracker, supports CSG, rectilinear array, and hexagonal array universe types.
4.1 Dynamic polymorphism (DP) method
This first method uses the standard multi-universe tracking algorithms put forth in Section 3.2, i.e., the approach used on the CPU in ORANGE. A tracker base class is created, defining the tracking operations listed in Table 1 as pure virtual methods. CSG and rectilinear array tracker classes inherit from the base class and implement the virtual functions. A polymorphic get_tracker function returns a pointer to either a CSG or rectilinear array tracker object. For a fair comparison, within CSG universes, the DP method uses BIH acceleration, which is described in Section 4.3.
4.2 Static polymorphism (SP) method
Like the DP method, the SP method uses separate tracker types for CSG and rectilinear array universes, but this is achieved with static polymorphism. Using this approach, each tracker operation involves a switch statement predicated on the type of the current universe (represented by an enumeration, UType). For example, a possible implementation of find_cell is shown in Listing 1. This function takes three arguments: the universe identifier (uid) of the current universe, the position, and a set of geometry parameters which supply a mapping between uid and UType. A CSG or rectilinear tracker is created based on the UType of the universe specified by the uid. The find_cell method is then called on the tracker, and the resulting cell is returned.
Listing 1: Possible C++ implementation of find_cell using static polymorphism. This code block has equivalent logic to the code in Listings 2, 3, and 4.
1__host__ __device__
2Cell find_cell(UniverseId uid,
3 Position pos,
4 Params params)
5{
6 switch(params.types[uid])
7 {
8 case UType::CSG:
9 auto tracker
10 = CSGTracker(uid, params)
11 return tracker.find_cell(pos);
12 case UType::RectArray:
13 auto tracker
14 = RectArrayTracker(uid,
15 params)
16 return tracker.find_cell(pos);
17 }
18}
In ORANGE, this behavior was achieved via a template metaprogramming approach which allows a single set of switch statement logic to be used by all tracking operations, as shown in Listings 2, 3, and 4. In Listing 2, a Traits struct is templated on UType, and template specializations are used to define the corresponding tracker type for each universe type. A visit_universe_type function returns the output of a given functor, which takes a Traits object as an argument. Listing 3 provides a visit_tracker function which returns the output of a supplied functor f, which takes a tracker of arbitrary type as an argument. This is achieved by wrapping f in a second functor—f2, which calls f for a given Traits object—and then calling visit_universe_type with this second functor.
Listing 4 shows how the visit_tracker function can be used to achieve polymorphism. A find_cell functor is first created. This functor, along with a uid and params, is passed visit_tracker, and the resulting cell is returned. Upon compilation, code in Listings 2, 3, and 4 produce bitcode equivalent to that of Listing 1. The other Table 1 tracking operations are implemented in an identical fashion to the find_cell example in Listing 4. As was the case with the DP method, CSG universes in the SP method also use BIH acceleration, which is described in Section 4.3.
Listing 2: Simplified C++ function for visiting different universe types.
1template<UType U>
2struct Traits;
3template<>
4struct Traits<UType::CSG>
5{
6 using tracker_type = CSGTracker;
7}
8template<>
9struct Traits<UType::RectArray>
10{
11 using tracker_type = RectArrayTracker;
12}
13
14template<class F>
15inline constexpr
16__host__ __device__ decltype(auto)
17visit_universe_type(F&& f, UType ut)
18{
19 switch (ut)
20 {
21 case UType::CSG:
22 return f(Traits<UType::CSG>{})
23 case UType::RectArray:
24 return f(Traits<UType::RectArray>{})
25 }
26}
Listing 3: Simplified C++ function for visiting different tracker types.
1template<class F>
2 __host__ __device__ decltype(auto)
3visit_tracker(F&& f,
4 UniverseId uid,
5 Params params)
6{
7 auto f2 = [&](auto traits){
8 return f(traits::tracker_type(
9 uid,
10 params));
11 }
12
13 return visit_universe_type(
14 f2,
15 params.universe_types[uid]);
16}
Listing 4: C++ implementation of find_cell using the visit_tracker function shown in Listing 3.
1auto find_cell = [&pos](auto&& tracker){
2 return tracker.find_cell(pos);
3 }
4Cell cell = visit_tracker(find_cell,
5 uid,
6 params)
4.3 Single tracker (ST) method
Instead of employing polymorphism, the final approach uses a single CSG tracker for CSG, rectilinear array, and hexagonal array universes. This is accomplished by converting rectilinear and hexagonal array universes into CSG universes. To do so, array cells are explicitly modeled as CSG cells using the method shown in Figure 2. This conversion is done automatically in ORANGE, and the resulting CSG universes are referred to as “pseudo-array universes,” a term specific to this work. Modeling arrays in this fashion is not a new approach: it is the simplest way of modeling an array within an MC code without specialized array universe types.
Because pseudo-array universes are CSG universes, they use the CSG tracking algorithms shown in Section 3.1 rather than the array versions that have improved time complexity. In other words, for pseudo-arrays, find_cell involves an $O(N)$ linear search over all cells and cross_surface requires an $O(N)$ linear search over all neighbor cells. The cross_surface algorithm has an additional complication for pseudo-array universes. In ORANGE, coincident surfaces are always deduplicated to simplify tracking, resulting in a large number of neighbor cells for pseudo-array surfaces. An example of this is shown in Figure 4. In this example, the magenta surface bounds all 32 colored cells. When crossing the magenta surface from the purple cell, the neighbor list includes all 31 gold cells.
Figure 4: Example of a pseudo-rectilinear-array neighbor list. A neutron in the purple cell crossing the magenta surface has all 31 gold cells as neighbors.
To improve the performance of tracking operations for CSG universes, including pseudo-array universes, a BIH acceleration structure was implemented. The BIH construction process is demonstrated in Figure 5. Bounding boxes for each cell are first ascertained. A partition plane is then chosen, and bounding boxes are partitioned into two sets according to the location of each bounding box center. For each of the two sets of partitioned bounding boxes, bounding planes are created by moving the partition to fully enclose all bounding boxes. By performing this process recursively on each set of bounding boxes, a binary tree structure is created, with edges specifying the half-spaces which contain all children.
Figure 5: BIH construction process demonstrated in 2D using the states of the American West. Map outline from OpenClipart (2023).
One key feature of BIH trees is that for a given node, the half spaces created by the bounding planes may overlap. This guarantees that each cell appears in the BIH tree exactly once, unlike $k$-D trees that must store a cell twice if its bounding box is bisected by a partition. This provides an advantage in the pseudo-array use case in which adjacent cells may have bounding boxes that overlap slightly because of floating point error. BIH trees can handle this case without significant performance consequences. BVH trees can also handle this overlap case. However, BVH trees may require more memory because they store six planes per node, significantly more than the two planes required by BIH trees. An open research question within the BIH construction process is how to choose partitions. For the purposes of this work, a standard surface area heuristic (SAH) partitioning scheme (Wald, 2007) was used. This was implemented by evaluating three equally spaced candidate partitions per axes at each partitioning step. Candidate partitions were evaluated using a cost function balancing the number of bounding boxes and the total surfaces area of bounding boxes appearing on each side of the partition.
The find_cell and cross_surface functions are accelerated by traversing the BIH tree. For find_cell, traversal is terminated when a cell is found that contains the supplied point. For cross_surface, the traversal is terminated when a cell is found that contains the supplied point, excluding the cell in which the neutron originated. When traversing the BIH tree, both edge conditions must be tested at each node because the half spaces are allowed to overlap. BIH trees are enabled for all CSG universes—not just pseudo-array universes. As mentioned in Sections 4.1 and 4.2, BIH trees are enabled for all CSG universes in the DP and SP methods as well.
5 Hardware description
All simulations were performed on the Summit and Frontier supercomputers at ORNL. Summit has 4608 compute nodes, each consisting of two 22-core IBM Power9 CPUs, and six NVIDIA Tesla V100 GPUs, each consisting of a single graphics complex die (GCD). One core per Power9 is reserved for system tasks, leaving 42 usable CPU cores per node. On the CPU side, 512 GB of RAM are available, and each V100 has 16 GB of RAM. Within this work, code was compiled on Summit with CUDA 11.5.2.
Frontier has 9408 compute nodes, each consisting of a 64-core AMD $\text{3}^{\text{rd}}$ Gen EPYC CPU, and four AMD Radeon Instinct MI250X, each consisting of a two GCDs. Eight CPU cores are reserved for system tasks, leaving 56 usable CPU cores per node. On the CPU side, Frontier also has 512 GB of RAM, but each MI250X has 128 GB of RAM. As a result, Frontier has four times as much GPU RAM per GCD compared to Summit (64 GB vs. 16 GB). The results presented here were compiled with ROCm 5.6.0. It is noted that large variations in performance (up to 50%) where observed as a function of compiler version.
6 Performance testing
The performance of the three multi-universe GPU tracking methods described in Section 4 relative to RTK (described in Section 3.3) was assessed by obtaining timing results for a rectilinear-array-based reactor problem. A full-core model of the NuScale SMR (Smith, 2017), shown in Figure 6, was chosen for this purpose. The small size of the the NuScale design—about one-eighth the size of a typical PWR—is economically attractive because of its low capital cost (Black et al., 2019), and also permits detailed full-core MC analysis. Likewise, this problem served as the challenge problem for the ExaSMR project within the Exascale Computing Project, in which Shift was used for GPU-based coupled neutron transport / thermal hydraulics analysis (Merzari et al., 2023).
The NuScale design consists of 37 assemblies arranged in a rectilinear grid. Each assembly is a 17$\times$17 array of pins and contains uranium dioxide fuel with a 235U enrichment of either 1.6%, 2.4%, or 3.1% (by mass). The inner circle of 3.1% enriched fuel assemblies contain borosilicate glass burnable neutron absorber rods. Spacer grids and nozzles have been homogenized into slabs for simplicity. For this analysis, the fresh (i.e., non-depleted) fuel version of the problem was used. Fresh fuel contains many fewer nuclides than depleted fuel, thereby minimizing time required to calculate cross sections and maximizing the relative time spent on tracking operations. As a result, inactive cycles with this problem (where no time is spent on tallies) represent the scenario in which tracking operations are expected to comprise the largest fraction of runtime.
Figure 6: Midplane radial (top) and axial (bottom) slices of a full core model of the NuScale SMR from Smith (2017).
For each of the four tracking methods, performance testing was conducted by measuring the tracking rate, i.e., the number of histories simulated per unit time, over a sweep of workloads. Here, workload refers to the number of histories per cycle assigned to each GPU GCD. Each trial consisted of 10 inactive cycles and 10 active cycles using all available GPU GCDs on a single node of Summit or Frontier. During active cycles, the neutron flux was was tallied on a $119\times 119\times 30$ superimposed rectilinear mesh (425,830 mesh volume elements).
Figure 7: Neutron tracking rates as a function of the number of histories run by each GPU GCD on Summit and Frontier for the NuScale SMR problem.
Tracking rate results are shown in Figure 7, noting that both plots have different $x$ scales but the same $y$ scale. As expected, the tracking rate increased with workload for all four methods, because large workloads allow the GPU to more effectively hide the latency of memory fetches associated with cross section calculations. Inactive cycle tracking rates are higher because no time is spent processing tally results. On Frontier, tracking rates for all four methods are near their asymptotic limit at a workload of $2.8\times 10^{7}$ histories/cycle/GCD, well before running out of memory, which was found to occur at $5\times 10^{7}$ histories/cycle/GCD for RTK and $3.7\times 10^{7}$ histories/cycle/GCD for the DP, SP, and ST methods. However, on Summit, which has 4×4$\times$4 × less RAM/GCD, all four methods ran out of RAM prior to nearing an asymptotic limit, with RTK running out memory at $1.2\times 10^{7}$ histories/cycle/GCD and the DP, SP, and ST methods running out of memory at $8\times 10^{6}$ histories/cycle/GCD.
Aside from the $1\times 10^{6}$ histories/s/GCD inactive cycles, in which the low workloads cause the kernel launch overhead to wash out any differences between the methods, the relationships between the performance of the four methods remained constant over all workloads. The highest tracking rates were achieved with RTK, as expected, followed by the SP method, the ST method, and the DP method. Further analysis was performed with the $8\times 10^{6}$ histories/cycle/GCD trials on Summit and $2.8\times 10^{7}$ histories/cycle/GCD trials on Frontier. Tracking rates for each of the four methods are shown in Table 2. Results show that all three experimental methods achieve over 90% of the RTK tracking rate on Summit. On Frontier, the DP, SP, and ST methods achieve 77.8%, 91.2%, and 83.4% of the tracking rate during inactive cycles, with all methods achieving at least 87.9% of the RTK tracking rate during active cycles. Therefore, whereas the SP method consistently provides the best performance, no method incurred a significant performance penalty.
Table 2: GPU tracking rates for the NuScale SMR problem, run on Summit with $8\times 10^{6}$ histories/cycle/GCD and Frontier with $2.8\times 10^{7}$ histories/cycle/GCD.
Summit Frontier
Method GPU tracking Fraction of GPU tracking Fraction of
rate RTK tracking rate RTK tracking
(neutron/s/GCD) rate (%) (neutron/s/GCD) rate (%)
Inactive RTK $1.38\times 10^{5}$ 100 $1.57\times 10^{5}$ 100
DP $1.27\times 10^{5}$ 92.2 $1.22\times 10^{5}$ 77.8
SP $1.36\times 10^{5}$ 98.5 $1.43\times 10^{5}$ 91.2
ST $1.28\times 10^{5}$ 93.3 $1.31\times 10^{5}$ 83.4
Active RTK $6.79\times 10^{4}$ 100 $7.62\times 10^{4}$ 100
DP $6.38\times 10^{4}$ 93.9 $6.69\times 10^{4}$ 87.9
SP $6.69\times 10^{4}$ 98.6 $7.26\times 10^{4}$ 95.3
ST $6.49\times 10^{4}$ 95.5 $6.93\times 10^{4}$ 91.0
The percentages of the total GPU runtime (inactive and active cycles) spent in tracking kernels are shown in Table 3. These values vary from 12.1–26.1%, indicating that tracking operations do not dominate runtime, as expected. Likewise, much larger differences in the performance of the four methods are observed when only considering the time spent within tracking kernels. Figure 8 shows the total time spent within the five principle tracking kernels introduced in Table 2. For a fair comparison, both Summit and Frontier results are from workloads of $8\times 10^{6}$ histories/cycle/GCD. This figure shows that the DP, SP, and ST methods spent 1.70$\times$, 1.14$\times$, and 1.50$\times$ more time in tracking kernels compared to RTK on Summit, and 2.45$\times$, 1.5$\times$, and 2.02$\times$ more time compared to RTK on Frontier. These results highlight the supremacy of the SP method, almost achieving parity with RTK on Summit.
Table 3: Fraction of total GPU runtime (inactive and active cycles) spent in geometry tracking kernels for the NuScale SMR problem. Summit results are for $8\times 10^{6}$ histories/cycle/GCD and Frontier results are for $2.8\times 10^{7}$ histories/cycle/GCD.
Runtime fraction (%)
Method Summit Frontier
RTK 12.1 13.3
DP 19.2 26.1
SP 13.6 18.3
ST 17.2 23.3
Figure 8: Total time spent in the 5 principle tracking kernels on Summit and Frontier for the NuScale SMR problem, both with $8\times 10^{6}$ histories/cycle/GCD. All tracking kernels not listed here accounted for less than 0.3% of the total tracking time.
Whereas overall tracking rates were slightly higher on Frontier, significantly more time was spent in tracking kernels on Frontier compared to Summit. Frontier also lags behind Summit when comparing GPU performance to CPU performance. Table 4 compares GPU to CPU tracking rates in terms of (1) the number of CPU cores required to match a single GPU GCD, and (2) the speedup, defined as the ratio of the total GPU tracking rate to the total CPU tracking rate on a single node. Noting that both GPU and CPU resources differ between these two machines, Summit has higher CPU core equivalence and speedups compared to Frontier in all cases. This can be attributed to the fact that significantly better CPU performance is achieved on Frontier, where a tracking rate of $4.94\times 10^{3}$ neutron/s/core was achieved for inactive cycles, compared to $2.82\times 10^{3}$ neutrons/s/core for inactive cycles on Summit.
Table 4: GPU to CPU performance comparison for the NuScale SMR problem, run on Summit with $8\times 10^{6}$ histories/cycle/GCD and Frontier with $2.8\times 10^{7}$ histories/cycle/GCD. Speedup refers to the reduction in time-to-solution from using all GPUs on a node compared to using all CPUs.
Summit Frontier
Method CPU core Speedup CPU core Speedup
equivs. equivs.
per GCD per GCD
Inactive RTK 48.9 6.98 31.7 4.53
DP 45.0 6.43 24.7 3.53
SP 48.1 6.87 28.9 4.13
ST 45.6 6.51 26.4 3.78
Active RTK 28.1 4.01 15.3 2.19
DP 26.3 3.76 13.4 1.92
SP 27.7 3.95 14.6 2.08
ST 26.8 3.83 13.9 1.99
Single node performance testing trials, which were run with only 10 inactive and 10 active cycles, were not sufficient to produce converged $k_{\mathrm{eff}}$ results. To confirm that all four tracking methods produce the same solution, a final trial was performed for each method on 50 nodes of Frontier with a workload of $2.5\times 10^{5}$ histories/cycle/GCD, for a total of $10^{8}$ histories per cycle. For each trial, 200 inactive cycles and 400 active cycles were run. A lower workload was necessary to complete all 600 cycles within Frontier’s walltime limit. Converged $k_{\mathrm{eff}}$ values agreed closely, as shown in Table 5. Neutron flux results for all 4 methods also matched expectations. Figure 9 shows the converged flux for the SP method trial. This figure clearly shows the depression on the flux within the borosilicate glass rods, which act as neutron absorbers.
[…]
Method $k_{\mathrm{eff}}$
RTK 1.046076 $\pm$ 0.000005
DP 1.046076 $\pm$ 0.000005
SP 1.046067 $\pm$ 0.000005
ST 1.046072 $\pm$ 0.000005
[…]
7 Demonstration problem
In Section 6, the SP method consistently provided the best performance, but all three methods achieved tracking rates reasonably close to that of RTK. From a software engineering perspective, when considering multiple approaches, the trade-off between performance and other factors such as code simplicity and maintainability must be assessed. The ST method achieves over 90% of the SP method tracking rate in all cases shown in Table 2 and has the added benefit of not requiring an additional hexagonal array tracker. This is advantageous for hexagonal-array-based reactors, which cannot be represented using RTK, and would otherwise require the development of a GPU-based hexagonal array tracker for use with the SP method. To demonstrate this benefit, the ST method was used to perform a $k_{\mathrm{eff}}$ calculation on the Empire microreactor benchmark problem, which consists of nested hexagonal arrays as shown in Figure 10. Microreactors, even smaller than SMRs, are designed to be factory assembled and easily transported in order to supply power for remote areas, disaster relief, and space applications.
The Empire microreactor consists of 18 hexagonal assemblies arranged around a central void region. Each assembly contains a total of 217 pins: 60 uranium nitride fuel pins with 16.05% 235U enrichment (by mass), 96 yttrium hydride moderator pins, and 61 stainless steel heat pipes filled with liquid sodium. Within this model, material within the heat pipe pins is homogenized for simplicity. The core is surrounded by 12 control drums which contain a europium boride neutron absorber on one side. These drums can be rotated to change the absorber configuration to control the reactivity within the core. For this analysis, the “drums-in” configuration was used, with all absorbers directly facing the core.
Figure 10: Midplane radial (top) and axial (bottom) slices of a full core model of the Empire microreactor in “drums-in” configuration.
The problem was run on 100 nodes on Frontier using all eight GCD s on each node, with a workload of $10^{6}$ histories/cycle/GCD, for a total of $8\times 10^{8}$ histories/cycle. To achieve converged results, 120 inactive cycles and 120 active cycles where performed. The flux and fission source were tallied on an extremely fine rectilinear mesh with 3350 mesh elements in $x$ and $y$ (0.05 cm resolution) and a single mesh element in $z$ spanning $\pm$0.25 cm on the midplane (11,222,500 mesh volume elements total). A converged $k_{\mathrm{eff}}$ of 1.026777 $\pm$ 0.000004 was obtained. Converged flux results are shown in Figure 11. This figure shows the depression in the flux around the neutron absorbers and interstitial spaces between assemblies where no fuel is present. Figure 12 shows the converged fission source, which is highly resolved within each fuel rod. The highest fission source intensity occurs around the edges of fuel rods. For rods in the outer parts of the assembly, the fission source is observed to be highest on the inward-facing edge.
Figure 11: Converged neutron flux distribution on the midplane of the Empire problem using the ST method, obtained on 100 nodes of Frontier with $10^{6}$ histories/cycle/GCD, with 120 inactive and 240 active cycles. Statistical errors in the flux on this slice are all less than 0.7% within the core region.
[…]
Table 6 shows tracking rates, as well as GPU vs. CPU comparison results, in the same format as that used in Table 4. For inactive cycles, the tracking rate of $5.29\times 10^{4}$ histories/s/GCD is 40.4% of the ST method tracking rate for the NuScale problem on Frontier. For active cycles, the tracking rate of $1.02\times 10^{4}$ histories/s/GCD is only 14.7% of the ST method tracking rate for the NuScale problem on Frontier. This is reasonable because the mesh used for the Empire problem is 26.4$\times$ larger than the mesh used for the NuScale problem and is much larger than meshes used for most practical applications. With respect to CPU execution, a modest 1.72$\times$ speedup was achieved for inactive cycles; however, for active cycles, a 2.94$\times$ slowdown was observed. The stark difference between inactive and active cycle performance can be attributed to tallying operations, which only occur during active cycles. Tally operations accounted for 61% of the total GPU runtime (i.e., runtime spend in both inactive and active cycles) as shown in Figure 13. These results indicate that more work is necessary to improve the performance of tally operations on the GPU.
Table 6: GPU to CPU performance comparison for the Empire problem on Frontier, with $10^{6}$ histories/cycle/GCD. Speedup refers to the reduction in time-to-solution from using all GPUs on a node compared to all CPUs.
Cycle GPU tracking CPU core Speedup
type rate equivs.
(neutron/s/GCD) per GPU
Inactive $5.29\times 10^{4}$ 12.0 1.72
Active $1.02\times 10^{4}$ 2.37 0.34
Figure 13: Percentage of total GPU runtime (both inactive and active cycles) spent performing geometry, physics, tally, and other operations for the Empire problem on Frontier. Cross section calculations (which typically dominate runtime) fall within the “physics” category.
8 Conclusion
In this work, three methods for GPU-based neutron tracking within multi-universe geometries where tested and compared to RTK. For the NuScale problem, the DP, SP, and ST methods spent 2.45$\times$, 1.50$\times$, and 2.02$\times$ more time in tracking kernels in inactive cycles on Frontier compared to RTK. However, although inactive cycles on this fresh fuel problem represent the maximum time spent performing tracking operations, these methods still achieved 77.8%, 91.2%, and 83.4% of the RTK tracking rate. It is concluded that any of these three methods—which unlike RTK can handle arbitrarily nested configurations—are suitable for production-level use, incurring only a modest performance penalty.
The SP method outperformed the DP and ST methods in nearly all cases. However, the ST method achieved over 90% of the SP method’s tracking rate in all cases, and unlike the SP method, it requires only a single CSG tracker. The flexibility of this approach was demonstrated with the Empire microreactor, a hexagonal array-based reactor which cannot be represented using RTK, and would require an additional GPU-based hexagonal array tracker to be written in order to use the SP method. The ST method was used to produce converged $k_{\mathrm{eff}}$, neutron flux, and fission source results, which matched expectations. A modest speedup of 1.72$\times$ over CPU execution on Frontier was observed during inactive cycles. Additional work is necessary outside the neutron tracking portions of Shift to remediate the a 2.94$\times$ slowdown observed for active cycles, understanding that the mesh tally used for this problem is far larger than mesh tallies used for typical applications. This work will facilitate GPU-based MC transport for reactor problems with geometry complexity beyond rectilinear arrays, as well as other non-reactor radiation transport problems with complex nested geometries.
Funding
This research was supported by the Exascale Computing Project (ECP), project number 17-SC-20-SC. The ECP is a collaborative effort of two DOE organizations, the Office of Science and the National Nuclear Security Administration, that are responsible for the planning and preparation of a capable exascale ecosystem—including software, applications, hardware, advanced system engineering, and early testbed platforms—to support the nation’s exascale computing imperative.
Acknowledgements
The authors would like to thank Stuart Slattery and Friederike Bostelmann for their internal technical reviews.